4. x + 4x + 13x = te-t, x(0) = 0, x(0) = 2.

use laplace transform
Answer #1


st + 4 x + 13x = et et a(o)=0 alo) - 2 Take cablau transform S?XIS) - 2 + 4SX1S) +13.X(S)= (5+12 = (s? + 45 +13) X(5) 2 t (